The probability that a patient recovers from a rare disease is 0.6. Calculate the
probability that out of 5 patients suffering from this disease, at least two would
recover.
1
Expert's answer
2019-04-28T17:17:05-0400

It is Binomial distribution with "p=0.6, n=5."


"P(X\geq2)=1-(P(X=0)+P(X=1))="

"=1-\begin{pmatrix}
 5 \\
 0
\end{pmatrix}0.6^0(1-0.6)^{5-0}-\begin{pmatrix}
 5 \\
 1
\end{pmatrix}0.6^1(1-0.6)^{5-1}="

"=1-0.01024-0.0768=0.91296"

"P(X\geq2)=0.91296"